# AdaptiveK Sparse Autoencoders: Dynamic Sparsity Allocation for Interpretable LLM Representations

### v1

@@ C.2 Other Evaluation Metrics
In this section, we evaluate the AdaptiveK SAE using five metrics from SAEBench [Karvonen et al., 2025]. For the Feature Absorption metric, we directly compare the results of AdaptiveK SAE with those of the baseline SAEs reported in SAEBench, noting that their training dataset was 2000 times larger than ours. For the remaining metrics, (Spurious Correlation Removal, Targeted Probe Perturbation, Resolving Attribute-Value Entanglements in Language Models, and Sparse Probing), the AdaptiveK SAE is compared against baseline SAEs trained on an identical amount of training data.
C.2.1 Feature Absorption
@@
Following [Karvonen et al., 2025], Feature Absorption is measured using a first-letter classification task. It first establishes a “ground truth” directional vector $p$, for each first-letter concept by training linear probes on the base language model’s activations ($a_{model}$). It then identifies a set of “main” SAE latents ($S_{main}$) expected to represent each letter. The core of the measurement involves analyzing individual instances (words). For an instance, if the main latents don’t fully capture the ground truth signal (i.e., $\sum_{i\in S_{\text{main }}}a_{i}d_{i}\cdot p<a_{\text{model }}\cdot p$, where $a_{i}$ a is latent activation and $d_{i}$ its decoder vector), and other “absorbing” latents ($S_{abs}$) that align with $p$ compensate for this deficit, an absorption fraction is calculated. Let $P_{main}=\sum_{j\in S_{\text{main }}}a_{j}d_{j}\cdot p$ be the projection from main features, and $P_{compensated\_by\_absorbers}$ be projection from the top few absorbing latents that cover the signal portion not captured by $P_{main}$. The instance-level absorption fraction $f_{abs}$ is then:
@@
Utilizing this metric, we assessed the performance of our AdaptiveK SAE on layer 3 of Pythia-160M and layer 12 of Gemma-2-2B, as shown in Figure 8 and 9. Notably, when directly benchmarked against SAEs from SAEBench [Karvonen et al., 2025] for Gemma-2-2B’s layer 12, AdaptiveK exhibited superior results on both metrics (Figure 10). This outperformance is particularly significant given that our AdaptiveK was trained on only 250,000 tokens, a dataset 2000 times smaller than the 500,000,000 tokens used for the SAEBench models. In AdaptiveK, concepts are correctly allocated to their intended primary latents rather than being dispersed across unrelated variables. Cases where concepts are entirely misrepresented in non-primary latents are notably rare. This demonstrates AdaptiveK’s exceptional effectiveness in maintaining conceptual integrity and preventing feature fragmentation.
@@
C.2.2 Spurious Correlation Removal (SCR)
Spurious Correlation Removal evaluates an SAE’s ability to disentangle distinct concepts by measuring how effectively it can remove spurious correlations from classifiers. Likewise, utilizing method from SAEBench [Karvonen et al., 2025], we first generated biased datasets containing spurious correlations (e.g., professor+male and nurse+female from the Bias in Bios dataset). A linear classifier is then trained on this biased dataset, learning to rely on both the target concept (profession) and the spurious concept (gender). To evaluate an SAE, the method identifies latents most strongly associated with the spurious concept (gender) through probe attribution scores. These identified latents are then zero-ablated, creating a modified classifier. The final SCR score is normalized as:
@@ C.2.2 Spurious Correlation Removal (SCR)
(a) SCR score with 10-latent ablation
(b) SCR score with 20-latent ablation
@@
C.2.3 Targeted Probe Perturbation (TPP)
@@
(a) TPP score with 10-latent ablation
(b) TPP score with 20-latent ablation
@@
C.2.4 Resolving Attribute-Value Entanglements in Language Models (RAVEL)
@@
In the disentanglement score (Figure 13(a)), AdaptiveK achieves 0.62, substantially higher than contemporary architectures like TopK and Matryoshka at comparable sparsity levels. For the cause score (Figure 13(b)), AdaptiveK reaches 0.6, roughly double the effectiveness of other SAEs at similar sparsity. This indicates AdaptiveK is exceptionally good at identifying and modifying the specific latents that control target attributes (like a city’s country). The isolation score (Figure 13(c)) shows AdaptiveK at 0.65, demonstrating it maintains unrelated attributes more effectively than others. While P Anneal and some other SAEs eventually reach similar or higher disentanglement scores, they require much higher sparsity levels (L0 > 1000) to do so, making them less practical for interpretability work that benefits from more compact representations. Combined with the earlier SCR and TPP results, these RAVEL findings confirm that AdaptiveK creates a latent space with exceptionally clean separation between different concepts and attributes, enabling more precise and controlled interventions on language model knowledge.
@@ C.2.5 Sparse Probing
Metric “Full Activations Accuracy” represents classification performance when using all SAE activations, establishing an upper bound. Metrics “Top-K Accuracy” (where K is 1, 2, or 5) measure performance when restricting the probe to only the K most relevant latents for each task. First, we measure information retention rate (SAE Full Activations Accuracy / LLM Full Activations Accuracy), which quantifies how well each SAE preserves the original model’s information when using all reconstructed activations. As shown in Figure 14, AdaptiveK’s retention ratio of approximately 0.997 demonstrates near-perfect preservation of the original model’s information. While Matryoshka’s slightly higher ratio (>1) suggests beneficial feature reorganization or denoising during reconstruction, such enhancement represents a supplementary advantage rather than a necessity.
@@
Appendix D Limitations and Broader Impacts
D.1 Limitations
@@ D.2 Broader Impacts
Prompt for Scoring Context Complexity Detailed Evaluation Dimensions 1. Lexical Complexity (Weight: 20%): Evaluate the vocabulary sophistication level using the following criteria Word Frequency: Proportion of uncommon words (not in the 5000 most frequent words) Word Length: Average syllable count and character length of words Lexical Diversity: Type–token ratio (unique words divided by total words) Technical Terminology: Presence of specialized or domain-specific vocabulary Lexical Density: Ratio of content words (nouns, verbs, adjectives, adverbs) to function words (pronouns, prepositions, articles, etc.) 2. Syntactic Complexity (Weight: 20%): Analyze sentence-structure complexity using these metrics Sentence Length: Average number of words per sentence Clause Density: Number of clauses per sentence Subordination: Frequency and depth of subordinate clauses Passive Voice: Proportion of sentences in passive voice Syntactic Variety: Diversity of sentence structures Embedding Depth: How deeply clauses are nested within one another 3. Conceptual Density (Weight: 25%): Assess the density and abstraction level of ideas presented Concept Count: Number of distinct concepts, ideas, or arguments introduced Concept Abstraction: Level of concreteness vs. abstraction of concepts Conceptual Networks: Complexity of relationships between concepts Information Density: Amount of information conveyed per paragraph Theoretical Complexity: Depth of theoretical constructs presented 4. Domain Specificity (Weight: 15%): Evaluate how much specialized domain knowledge is required Background Knowledge: Prerequisite knowledge assumed by the text Domain Vocabulary: Concentration of field-specific terminology Conceptual Familiarity: How familiar concepts would be to general readers Specialized References: References to domain-specific methods, theories, or figures Audience Specificity: How targeted the text is to specialists vs. general readers 5. Logical Structure (Weight: 10%): Analyze the complexity of reasoning patterns Argument Structure: Complexity of argumentative or explanatory structure Logical Operations: Presence of conditional, causal, comparative reasoning Inference Requirements: Extent to which the reader must infer rather than being told explicitly Logical Connections: Clarity and complexity of connections between ideas Reasoning Chains: Length and complexity of logical chains 6. Contextual Dependencies (Weight: 10%): Assess how much the text relies on external context Intertextual References: References to other texts or knowledge sources Cultural Knowledge: Required cultural or historical background Implicit Information: Amount of information that remains unstated yet necessary Presuppositions: Assumptions the text makes about reader knowledge Discourse Context: Degree to which meaning depends on broader discourse context Text to Evaluate {text} Required Output Format Only return a JSON object with the following structure:

### v2

@@ F.3 Other Evaluation Metrics
In this section, we evaluate the AdaptiveK SAE using five metrics from SAEBench (Karvonen et al., 2025). For the Feature Absorption metric, we directly compare the results of AdaptiveK SAE with those of the baseline SAEs reported in SAEBench, noting that their training dataset was 2000 times larger than ours. For the remaining metrics, (Spurious Correlation Removal, Targeted Probe Perturbation, Resolving Attribute-Value Entanglements in Language Models, and Sparse Probing), the AdaptiveK SAE is compared against baseline SAEs trained on an identical amount of training data.
F.3.1 Feature Absorption
@@
Following (Karvonen et al., 2025), Feature Absorption is measured using a first-letter classification task. It first establishes a “ground truth” directional vector $p$, for each first-letter concept by training linear probes on the base language model’s activations ($a_{model}$). It then identifies a set of “main” SAE latents ($S_{main}$) expected to represent each letter. The core of the measurement involves analyzing individual instances (words). For an instance, if the main latents don’t fully capture the ground truth signal (i.e., $\sum_{i\in S_{\text{main }}}a_{i}d_{i}\cdot p<a_{\text{model }}\cdot p$, where $a_{i}$ a is latent activation and $d_{i}$ its decoder vector), and other “absorbing” latents ($S_{abs}$) that align with $p$ compensate for this deficit, an absorption fraction is calculated. Let $P_{main}=\sum_{j\in S_{\text{main }}}a_{j}d_{j}\cdot p$ be the projection from main features, and $P_{compensated\_by\_absorbers}$ be projection from the top few absorbing latents that cover the signal portion not captured by $P_{main}$. The instance-level absorption fraction $f_{abs}$ is then:
@@
Utilizing this metric, we assessed the performance of our AdaptiveK SAE on layer 3 of Pythia-160M and layer 12 of Gemma-2-2B, as shown in Figure 19 and 20. Notably, when directly benchmarked against SAEs from SAEBench (Karvonen et al., 2025) for Gemma-2-2B’s layer 12, AdaptiveK exhibited superior results on both metrics (Figure 21). This outperformance is particularly significant given that our AdaptiveK was trained on only 250,000 tokens, a dataset 2000 times smaller than the 500,000,000 tokens used for the SAEBench models. In AdaptiveK, concepts are correctly allocated to their intended primary latents rather than being dispersed across unrelated variables. Cases where concepts are entirely misrepresented in non-primary latents are notably rare. This demonstrates AdaptiveK’s exceptional effectiveness in maintaining conceptual integrity and preventing feature fragmentation.
@@
F.3.2 Spurious Correlation Removal (SCR)
Spurious Correlation Removal evaluates an SAE’s ability to disentangle distinct concepts by measuring how effectively it can remove spurious correlations from classifiers. Likewise, utilizing method from SAEBench (Karvonen et al., 2025), we first generated biased datasets containing spurious correlations (e.g., professor+male and nurse+female from the Bias in Bios dataset). A linear classifier is then trained on this biased dataset, learning to rely on both the target concept (profession) and the spurious concept (gender). To evaluate an SAE, the method identifies latents most strongly associated with the spurious concept (gender) through probe attribution scores. These identified latents are then zero-ablated, creating a modified classifier. The final SCR score is normalized as:
@@ F.3.2 Spurious Correlation Removal (SCR)
(a) SCR with 10-latent ablation
(b) SCR with 20-latent ablation
@@
F.3.3 Targeted Probe Perturbation (TPP)
@@
(a) TPP with 10-latent ablation
(b) TPP with 20-latent ablation
@@
F.3.4 Resolving Attribute-Value Entanglements in Language Models (RAVEL)
@@
In the disentanglement score (Figure 24(a)), AdaptiveK achieves 0.62, substantially higher than contemporary architectures like TopK and Matryoshka at comparable sparsity levels. For the cause score (Figure 24(b)), AdaptiveK reaches 0.6, roughly double the effectiveness of other SAEs at similar sparsity. This indicates AdaptiveK is exceptionally good at identifying and modifying the specific latents that control target attributes (like a city’s country). The isolation score (Figure 24(c)) shows AdaptiveK at 0.65, demonstrating it maintains unrelated attributes more effectively than others. While P Anneal and some other SAEs eventually reach similar or higher disentanglement scores, they require much higher sparsity levels (L0 >> 1000) to do so, making them less practical for interpretability work that benefits from more compact representations. Combined with the earlier SCR and TPP results, these RAVEL findings confirm that AdaptiveK creates a latent space with exceptionally clean separation between different concepts and attributes, enabling more precise and controlled interventions on language model knowledge.
@@ F.3.5 Sparse Probing
Metric “Full Activations Accuracy” represents classification performance when using all SAE activations, establishing an upper bound. Metrics “Top-K Accuracy” (where K is 1, 2, or 5) measure performance when restricting the probe to only the K most relevant latents for each task. First, we measure information retention rate (SAE Full Activations Accuracy / LLM Full Activations Accuracy), which quantifies how well each SAE preserves the original model’s information when using all reconstructed activations. As shown in Figure 25, AdaptiveK’s retention ratio of approximately 0.997 demonstrates near-perfect preservation of the original model’s information. While Matryoshka’s slightly higher ratio (>>1) suggests beneficial feature reorganization or denoising during reconstruction, such enhancement represents a supplementary advantage rather than a necessity.
@@
F.4 Hyperparameters Analysis
F.4.1 Regularization strength $\lambda$
As stated in Section 3.1.1, to determine the optimal regularization strength $\lambda$, we perform 5-fold cross-validation. For each $\lambda$ in the set {0.001, 0.01, 0.1, 1.0, 10.0, 100.0, $\{0.001,\,0$.01,\,0.1,\,1.0,\,10.0,\,100.0,\,1000.0\}, the probe is trained on four folds and evaluated on the remaining fold using root mean squared error (RMSE). Table 4 reports the average RMSE across folds for each $\lambda$ on selected models, with the best $\lambda$ highlighted in bold.
Table 4: Performance across models with varying $\lambda$ values
$\lambda$ Gemma-2-2b Gemma-2-9b Llama-3.1-8b Qwen-3-8b Qwen-3-14b Phi-4-14b
0.001 1.1937 1.2663 1.2612 1.0875 1.2425 1.1934
0.01 1.1937 1.2663 1.2612 1.0875 1.2424 1.1934
0.1 1.1937 1.2663 1.2612 1.0875 1.2425 1.1934
1.0 1.1937 1.2662 1.2612 1.0875 1.2425 1.1934
10.0 1.1937 1.2663 1.2611 1.0874 1.2424 1.1934
100.0 1.1935 1.2663 1.2605 1.0873 1.2424 1.1933
1000.0 1.1937 1.2663 1.2608 1.0875 1.2424 1.1933
F.4.2 Steepness of sigmoid function $s$
$s$ determines how complexity scores map to k-values. Our experiments on Gemma-2-2B (Table 5) show that $s$ exhibits remarkable robustness across the tested range of 2.0 to 12.0. As $s$ increases, the k-value distribution becomes more dynamic: lower values (ss=2.0) produce conservative allocation with a narrow range (min k=143, max k=225), while higher values (ss=12.0) enable more aggressive feature allocation with expanded ranges (min k=58, max k=315). Crucially, reconstruction quality remains stable across all tested values. These results indicate that our method is highly robust to $s$ parameter selection, with performance variations of less than 1% across the entire range, validating our choice of ss=6.0 as a balanced default configuration.
Table 5: Effect of $s$ on $k$ statistics and performance metrics
$k$ on test set min $k$ max $k$ avg $k$ Explained Variance Cosine Similarity L2 Ratio
$s=2.0$ 143 225 188 0.738 0.908 0.911
$s=4.0$ 118 267 203 0.741 0.909 0.914
$s=6.0$ 96 291 214 0.743 0.909 0.921
$s=8.0$ 80 304 222 0.743 0.909 0.916
$s=12.0$ 58 315 232 0.742 0.909 0.917
F.4.3 Probe weight $\gamma$
$\gamma$ balances the SAE reconstruction loss and probe loss during joint fine-tuning, where smaller values prioritize reconstruction quality and produce more conservative k-value allocation, while larger $\gamma$ values enhance probe dominance and expand the k-value range to better reflect complexity differences. Our experiments (Table 6) confirm this expected behavior, with k-value ranges expanding as $\gamma$ increases (from min k=98, max k=285 at $\gamma$=0.5 to min k=92, max k=293 at $\gamma$=1.0), but across $\gamma\in[0.5,1.0]$ show minimal performance variation. This stability demonstrates that our method is robust to $\gamma$ selection, with performance variations under 1%, validating our default choice of $\gamma$=0.9.
Table 6: Effect of $\gamma$ on $k$ statistics and performance metrics
$k$ on test set min $k$ max $k$ avg $k$ Explained Variance Cosine Similarity L2 Ratio
$\gamma=0.5$ 98 285 208 0.742 0.909 0.913
$\gamma=0.7$ 97 289 210 0.743 0.909 0.916
$\gamma=0.9$ 96 291 214 0.743 0.909 0.921
$\gamma=1.0$ 92 293 215 0.743 0.909 0.919
F.4.4 Deviation penalty $\delta$
$\delta$ prevents probe parameters from deviating too far from their pre-trained values. Starting with $\delta$=0.2, we dynamically adjust this weight during training by monitoring probe loss changes over the recent 3 steps. We calculate the loss change rate as $\frac{\text{earliest loss}-\text{latest loss}}{\text{earliest loss}}$ and adjust $\delta$ accordingly: when loss decreases rapidly (change rate >> 0.05), indicating good probe learning progress in complexity prediction, we reduce the deviation constraint by setting $\delta=\text{current}_{\delta}\times 0.8\quad(\text{minimum }0.01)$ to allow more parameter flexibility; when loss stagnates or increases, suggesting learning difficulties or overfitting, we strengthen the constraint by setting $\delta=\text{current}_{\delta}\times 1.2\quad(\text{maximum }0.5)$ to prevent excessive parameter drift. The upper bound of 0.5 is critical because at this value, the deviation loss becomes comparable in magnitude to the probe loss, and higher values would make the deviation penalty too dominant, completely preventing the probe from adapting to new patterns when learning becomes difficult. Additionally, since deviation loss measures distance from initial values, excessive weights would create large gradients that harm training stability.
F.4.5 Regarding Sigmoid-based Transformation
The sigmoid function provides a smooth non-linear mapping that aligns well with empirical observations about complexity–feature relationships. Our implementation uses only two intuitive parameters: mid_point, which specifies the complexity level corresponding to the base $k$, and steepness, which controls the transition smoothness. In practice, we found these parameters to be stable across models. Setting mid_point=0.5 centers the allocation around normalized complexity, while steepness=6.0 ensures appropriate transition smoothness. To test robustness, we also experimented with Linear Mapping and Exponential Mapping, with results on Gemma-2-2B reported in Table 7. All three mapping methods achieved similar performance across metrics, demonstrating that our approach is not sensitive to the specific choice of mapping.
Table 7: Comparison of mapping methods
Method avg $k$ min $k$ max $k$ Explained Var Cosine Sim L2 Ratio Rel Recon Bias
Sigmoid 214 95 291 0.80 0.93 0.93 0.9995
Linear 206 98 296 0.79 0.93 0.92 0.9988
Exponential 144 53 271 0.78 0.92 0.92 0.9949
F.5 SAE Performance with Larger k Value
Additional experiments have been conducted with different $k_{min}$ and $k_{max}$. As shown in Table 8, when scaling $k_{max}$ from 320 to 640, we observe steady improvements: Explained Variance increases from 0.743 to 0.789, Cosine Similarity from 0.909 to 0.926, and L2 Ratio from 0.921 to 0.935. These results indicate that AdaptiveK benefits from increased capacity similarly to standard SAEs.
Table 8: Performance with different $k_{\min}$ and $k_{\max}$ settings
$k$ on test set min $k$ max $k$ avg $k$ Explained Variance Cosine Similarity L2 Ratio
$k_{\min}=20,\;k_{\max}=320$ 96 291 214 0.743 0.909 0.921
$k_{\min}=20,\;k_{\max}=480$ 132 435 313 0.768 0.919 0.926
$k_{\min}=20,\;k_{\max}=640$ 170 579 415 0.789 0.926 0.935
Appendix G Training Efficiency Analysis
In large-scale or token-level training scenarios, complexity annotation is often regarded as a scalability challenge, since exhaustive annotation may become costly. AdaptiveK addresses this issue through several design considerations.
Firstly, annotation is a separate process. Complexity annotation occurs during the data preprocessing stage as targets for complexity prediction, independent of SAE’s three-stage training. Once completed, the annotations can be used to train linear probes for any model architecture. Secondly, for large-scale training data, only a subset is annotated for linear probe training while using the full dataset for SAE training.
Training efficiency is further assessed through comparisons across multiple SAE configurations. While traditional SAEs avoid complexity annotation, they require training multiple SAEs with different sparsity settings (different k values or sparsity penalties). Table 9 shows complete training times for other SAEs with single sparsity configurations on Gemma-2-2B, all exceeding AdaptiveK’s training time. For six different sparsity levels (e.g., k=20,40,80,160,320,640), the total training time would exceed AdaptiveK by more than 6-fold.
Table 9: Total training time (minutes) for different SAEs
AdaptiveK Batch TopK Gated Matryoshka P Anneal Relu Relu New TopK
Total (min) 11084 13853 13908 13773 13902 13835 13814 13955
Appendix H Adaptability to Token-Level Evaluation
Although AdaptiveK operates at the context level by default, its design also enables reliable token-level evaluation. The training process uses contexts of 1024 tokens each (as mentioned in Section 3.1.1). These contexts contain multiple sentences of varying lengths, including truncated incomplete sentence. The complexity of the last token in each context is used to represent the entire context, ensuring generalization capability. This deliberate approach of not using complete sentences enables training with the last token regardless of sentence length (whether 32 tokens, 256 tokens, etc.).
Due to training process, during evaluation, even when the input context is not 1024 tokens (for example, the first 256 or 500 tokens), our method can still effectively compute complexity based on the representation of the last token in the first n tokens. This is the reason and mechanism for why and how it can effectively predict the complexity of tokens at arbitrary positions.
Empirical results on Gemma-2-9B further confirm this adaptability. As shown in Table 10 reconstruction performance is consistent across token positions 10, 100, 200, 500, 800, and 1000. The cosine similarity varies by only 0.3% (0.954-0.957), indicating highly consistent reconstruction directions. L2 ratios and reconstruction bias remain close to 1.0 (0.9583-0.9701, 1.0052-1.0208), demonstrating accurate reconstruction magnitudes. All metrics vary minimally (<< 5%) across positions, proving that AdaptiveK SAE maintains stable reconstruction quality at any context position. Our context-level approach reflects efficiency considerations rather than methodological limitations.
Table 10: Evaluation across different token positions
Token Position Cosine Similarity Explained Variance L2 Ratio Recon Bias K Value
10 0.9544 0.7865 0.9701 1.0208 267
100 0.9570 0.8255 0.9688 1.0130 282
200 0.9557 0.8268 0.9622 1.0104 285
500 0.9544 0.8203 0.9674 1.0130 285
800 0.9557 0.8333 0.9622 1.0104 286
1000 0.9557 0.8151 0.9583 1.0052 290
Appendix I Limitations and Broader Impacts
I.1 Limitations
@@ I.2 Broader Impacts
Prompt for Scoring Context Complexity Detailed Evaluation Dimensions 1. Lexical Complexity (Weight: 20%): Evaluate the vocabulary sophistication level using the following criteria Word Frequency: Proportion of uncommon words (not in the 5000 most frequent words) Word Length: Average syllable count and character length of words Lexical Diversity: Type–token ratio (unique words divided by total words) Technical Terminology: Presence of specialized or domain‑specific vocabulary Lexical Density: Ratio of content words (nouns, verbs, adjectives, adverbs) to function words (pronouns, prepositions, articles, etc.) 2. Syntactic Complexity (Weight: 20%): Analyze sentence‑structure complexity using these metrics Sentence Length: Average number of words per sentence Clause Density: Number of clauses per sentence Subordination: Frequency and depth of subordinate clauses Passive Voice: Proportion of sentences in passive voice Syntactic Variety: Diversity of sentence structures Embedding Depth: How deeply clauses are nested within one another 3. Conceptual Density (Weight: 25%): Assess the density and abstraction level of ideas presented Concept Count: Number of distinct concepts, ideas, or arguments introduced Concept Abstraction: Level of concreteness vs. abstraction of concepts Conceptual Networks: Complexity of relationships between concepts Information Density: Amount of information conveyed per paragraph Theoretical Complexity: Depth of theoretical constructs presented 4. Domain Specificity (Weight: 15%): Evaluate how much specialized domain knowledge is required Background Knowledge: Prerequisite knowledge assumed by the text Domain Vocabulary: Concentration of field‑specific terminology Conceptual Familiarity: How familiar concepts would be to general readers Specialized References: References to domain‑specific methods, theories, or figures Audience Specificity: How targeted the text is to specialists vs. general readers 5. Logical Structure (Weight: 10%): Analyze the complexity of reasoning patterns Argument Structure: Complexity of argumentative or explanatory structure Logical Operations: Presence of conditional, causal, comparative reasoning Inference Requirements: Extent to which the reader must infer rather than being told explicitly Logical Connections: Clarity and complexity of connections between ideas Reasoning Chains: Length and complexity of logical chains 6. Contextual Dependencies (Weight: 10%): Assess how much the text relies on external context Intertextual References: References to other texts or knowledge sources Cultural Knowledge: Required cultural or historical background Implicit Information: Amount of information that remains unstated yet necessary Presuppositions: Assumptions the text makes about reader knowledge Discourse Context: Degree to which meaning depends on broader discourse context Text to Evaluate {text} Required Output Format Only return a JSON object with the following structure:
@@
Appendix J The Use of Large Language Models (LLMs)
We utilized GPT-4.1-mini API to generate complexity scores for training contexts. Each 1024-token context was evaluated across six semantic dimensions: lexical complexity, syntactic complexity, conceptual density, domain specificity, logical structure, and contextual dependencies, producing normalized complexity scores from 0 to 10. Additionally, we employed LLMs for language polishing and grammatical refinement of the manuscript.